The city of a population of 161,565 has requested to evaluate its solids waste collection system.

Determine:

i) mean time per collection stop plus the mean time to reach the next stop,

ii) number of pickup locations per load

iii) minimum number of trucks the city must own.

Consider the following data:

a) Average truck capacity =18 m3

b) Average observed compaction ratio =3.28

c) Crew size = 2 persons

d) Number of pickups = 1/week (no rear-of-house service)

e) Average number of residence per stop= 2.5

f) Average uncompact density = 122 kg/m3

g) Average transport time to disposal site including delay and dumping = 1.5 h/trip

h) Average number of trips to disposal site =2/day

i) Rest breaks = 2 at 15 min

j) Average maintenance downtime = 36 min/day

k) Average work day = 8 hours

l) Average percentage of truck out of service for majority repairs = 15%

m) Average number of cans per stop = 2.5/wk at 0.0911 m3/can


1
Expert's answer
2022-03-24T01:41:05-0400

i.) mean time per collection stop plus the mean time to reach the next stop


Vp = (Average number of cans per stop) "\times" (Volume per can)


Vp = 2.95 "\times" 0.0911


Vp = 0.2689 m3/stop


ii.) number of pickup locations per load


Np = Number of pickups "\times" Average number of residence per stop "\times" Average number of trips to disposal site


Np = 1"\times" 2.5 "\times" 2


Np = 5


iii.) minimum number of trucks the city must own

Nt = Number of pickups"\times" Average number of residence per stop


Nt = "1\times2.5"


Nt = 3 trucks